# Multi-Level Embedding and Alignment Network with Consistency and Invariance Learning for Cross-View Geo-Localization

### v1

@@ IV-F Visualization
Furthermore, to further validate the performance of MEAN in multiscale environments and under significant viewpoint differences, we visualized the matching performance on the SEUS-200 dataset at various altitudes (150m, 200m, 250m, and 300m) as shown in Fig. 6. The results indicate that even in scenarios with substantial scale and perspective discrepancies, MEAN maintains more stable modality invariance between different viewpoints and effectively learns modality consistency across multiple height settings.
@@ V CONCLUSION
In this paper, we propose a lightweight multi-level embedding alignment network (MAEN) framework for cross-view geo-localization tasks. The framework aims to enhance feature representation capability and discriminability through a multi-branch structure, comprising progressive diversification embedding, global extension embedding, and cross-domain enhanced alignment. Specifically, the progressive diversification embedding branch focuses on generating diverse feature embeddings to accommodate complex geographic view variations, utilizing contrastive learning to improve feature consistency and discriminability across domains. The global extension embedding branch further optimizes the interaction between global and fine-grained features, enabling coherent expression of cross-domain information. The cross-domain enhanced alignment branch learns a shared mapping between domains through adaptive calibration and applies a cross-domain invariance alignment loss to overcome the limitations of relying on local detail alignment. This strengthens the intrinsic correlation of cross-domain features and further mines consistency and invariance within the embedding space. Experimental results demonstrate that MAEN achieves competitive performance, striking a notable balance between matching accuracy and computational efficiency, while exhibiting superior adaptability and robustness in cross-view geo-localization settings, outperforming existing state-of-the-art methods in certain cases.

### v2

@@ IV-G Retrieval Results
Furthermore, to further validate the performance of MEAN in multiscale environments and under significant viewpoint differences, we visualize the matching performance on the SEUS-200 dataset at various altitudes (150m, 200m, 250m, and 300m) as shown in Fig. 6. The results indicate that even in scenarios with substantial scale and perspective discrepancies, MEAN maintains more stable modality invariance between different viewpoints and effectively learns modality consistency across multiple height settings.
@@ V CONCLUSION
In this paper, we propose a lightweight MAEN framework for cross-view geo-localization tasks. The framework aims to enhance feature representation capability and discriminability through a multi-branch structure, comprising progressive diversification embedding, global extension embedding, and cross-domain enhanced alignment. The progressive diversification embedding branch focuses on generating diverse feature embeddings to accommodate complex geographic view variations and utilizing contrastive learning to improve feature consistency and discriminability. The global extension embedding branch further optimizes the interaction between global and fine-grained features to realize the coherent expression of cross-domain information. The cross-domain enhanced alignment branch learns a shared mapping between domains through adaptive calibration and applies a cross-domain invariance alignment loss to overcome the limitations of relying on local detail alignment. This strengthens the intrinsic correlation of cross-domain features and further mines consistency and invariance within the embedding space. Experimental results demonstrate that MAEN achieves competitive performance to strike a notable balance between matching accuracy and computational efficiency, while exhibiting superior adaptability and robustness in cross-view geo-localization tasks and outperforming existing state-of-the-art methods in certain cases.